OpenStudy (anonymous):

If two adults and three children go to the movies, it costs $31.50. If one adult and five children go to the movies, it costs $35.00. What is the cost of one adult movie ticket? $5.50 $6.50 $7.50

OpenStudy (acacia):

make an equation x=adult price y=children price first senario: 2 adults and 3 children => 2x+3y=31.50 second senario: 1 adult and 5 children => x+5y=35.00 then solve for x by elimination. -5(2x+3y=31.50) =\ -10x-15y=-157.5 3(x+5y=35.00) =/ 3x+15y=105 why multiply by these numbers? because you have to get one set of variables to cancel out each other when you add